OpenStudy (anonymous):

state the slope and y-intercept of y=0.12x-3.75

OpenStudy (anonymous):

the slope is m so it would be 3, and the y intercept is b, so it is 2

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

:) for those kind of problems use the formula y=mx+b its the easiest

OpenStudy (anonymous):

wait im sorry i was looking at something else.

OpenStudy (anonymous):

the first one said the equation was y=3x+2 but now its different. which one is it?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

ok well then hold on.

OpenStudy (anonymous):

is 3.75 negative?

OpenStudy (anonymous):

no its um y=0.12x minus 3.75

OpenStudy (anonymous):

the slope is 0.12 and the y intercept is -3.75 :)

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about this -y=6x+4

OpenStudy (anonymous):

6 is the slope, and 4 is the y-intercept

OpenStudy (anonymous):

what about 6y=3x-9

OpenStudy (anonymous):

3 is the slope, -9 is the y-intercept

OpenStudy (anonymous):

what about 2-5x=5y

OpenStudy (anonymous):

slope is -5, y-intercept is 2

OpenStudy (anonymous):

what about 6y+42=5x

OpenStudy (anonymous):

5 is the slope, 42 is the y-intercept

OpenStudy (anonymous):

what about this write an equation slope=5/8,passes through origin

OpenStudy (anonymous):

y=5x+8